OpenStudy (anonymous):

find a vector equation and parametric equations for the line through point (0,14,-10) and parallel to the line x=-1+2T, y=6-3T, Z=3+9T

OpenStudy (amistre64):

parallel have same vector parts; just different point parts

OpenStudy (amistre64):

x = -1+2t y = 6-3t z = 3+9t strip off the old point (-1,6,3) x = Px+2t y = Py-3t z = Pz+9t and add in the new point (0,14,-10) x = 0+2t y = 14-3t z = -10+9t and we get the parametric

OpenStudy (amistre64):

what a vector equation look like? is the that v = i+j+k stuff?

OpenStudy (anonymous):

what does the 2T, -3T and 9T represent?

OpenStudy (anonymous):

yes it is

OpenStudy (amistre64):

the vector <2,-3,9> strecthed to any scalar "t"

OpenStudy (amistre64):

to form a line; its just a vector attached to a pointand strecthched to hit all point on the line

OpenStudy (amistre64):

v = t(2i -3j +9) maybe?

OpenStudy (amistre64):

i got no idea how your text formats a vector equation, but its pretty basic from that i believe

OpenStudy (anonymous):

okay how would i do it if it was perpendicular instead of parallel?

OpenStudy (amistre64):

the perp vector has to dot with the original vector and get 0

OpenStudy (amistre64):

2,-3,9 a, b, c ------- 2a-3b+9c = 0

OpenStudy (anonymous):

okay still kinda confused, i'll ask another question on it hang on

OpenStudy (amistre64):

there are infinelty many perp vector to a line; they form a plane all around it

OpenStudy (amistre64):